# Federated Deconfounding and Debiasing Learning for Out-of-Distribution Generalization

### v1

@@ 1 Introduction
Federated out-of-distribution (OOD) generalization typically leverages data with diverse attributes across clients for collaborative modeling. The aim is to enhance model performance on unseen distributions Qi et al. (2024a); Liao et al. (2024). It enables federated learning (FL) clients to independently train models and engage in multiple iterations of parameter-level interactions with the FL server, which can optimize and aggregate models from different sources without exposing private data Kairouz et al. (2021); Wen et al. (2023). Despite its contributions to data privacy protection, existing studies show that attribute skew among sources often leads to performance degradation in FL models Abdelmoniem et al. (2023); Zhang et al. (2023); Yang et al. (2025). This is mainly because local models often establish spurious associations between specific attributes and labels, making it challenging to aggregate multiple local models with poor generalization ability into a strong global model.
@@
Existing approaches for enhancing FL model performance on OOD samples can be broadly divided into two groups: 1) knowledge distillation-based methods Wei and Han (2024); Luo et al. (2022); Yu et al. (2023); Wang et al. (2024) and 2) data augmentation-based methods Chen et al. (2023); Shenaj et al. (2023); Park et al. (2024). The former either treats the teacher’s knowledge (e.g., the output of global or other local models) as a regularizer to provide additional supervision or decouples category-irrelevant features in the latent space, thereby mitigating the interference of irrelevant attributes. For instance, MCGDM Wei and Han (2024) shares classifiers across sources to match gradients both within and across domains to reduce overfitting to attribute features. DFL Luo et al. (2022) applies mutual information-based disentanglement to address negative transfer due to attribute skew. However, the suboptimal performance of the teacher model on OOD samples and the single environment limit the effects of knowledge distillation and attribute decoupling. The latter approach typically trains generators or uses pre-trained generative models to create data with diverse attributes (e.g., CCST Chen et al. (2023)). For example, FedCCRL Wang and Tang (2024) fuses different samples for data augmentation. However, they often introduce privacy risks with performance limited by the quality of the generated data.
@@ 2.1 Data Augmentation-based Methods
Skewed attribute distributions across clients often cause models to overfit on domain-specific features, leading to amplified spurious relationships between these features and the labels. To address this, data augmentation-based methods Wang and Tang (2024); Shenaj et al. (2023); Liu and et al. (2021); Xu et al. (2023); Zhang et al. (2024); de Luca et al. (2022) have been developed to improve model generalization to previously unseen attributes by boosting the diversity of the data attributes. These methods typically rely on two approaches: 1) training data generators to create novel samples, or 2) utilizing pre-trained diffusion models to enrich sample diversity. The first approach involves exchanging local information between clients to generate new samples from different domains. For instance, FIST Nguyen et al. (2024) and StableFDG Park et al. (2024) produce samples with varying styles by sharing style-related information across clients. The second approach leverages prompt-driven techniques to generate samples that align with certain criteria Morafah et al. (2024); Zhao et al. (2023). Although enhancing diversity, these strategies often introduce privacy concerns. Moreover, the quality differences between generated and original samples also limit their effectiveness.
@@ 2.2 Knowledge Distillation-based Methods
Knowledge distillation-based approaches utilize generalized knowledge to guide local models in extracting shared features that are independent of specific attributes. These methods typically adopt two main strategies: 1) regularizing representations from different sources to align them, thus promoting consistency in representations across diverse contexts Wang et al. (2023); Qi et al. (2024b); Liu and et al. (2021); Sun et al. (2023); or 2) decoupling invariant features within the latent space to minimize the impact of confounding factors Yu et al. (2023). For instance, FedProc Mu et al. (2023) and FPL Huang and et al. (2023) leverage prototypes to enforce consistent representation learning across all clients, thereby encouraging them to converge within a shared representation space. MCGDM Wei and Han (2024) mitigates overfitting within individual domains by employing both intra-domain and cross-domain gradient matching to improve generalization. Similarly, FedIIR Guo and et al. (2023) enables the model to implicitly learn invariant relationships by capitalizing on prediction inconsistencies and gradient alignment across clients. Although these techniques have achieved notable performance improvements, they often rely on single-domain data, which hinders model transferability to OOD domains. Our FedDDL method bridges these gaps.
@@ 3 Preliminaries
An FL system typically involves multiple distinct data sources (a.k.a., FL clients), denoted as $\mathcal{S}=\{S_{1},S_{2},\dots,S_{K}\}$, and a central coordinating server $C$. Each $S_{k}$ utilizes its own private dataset $D^{k}=\{(X^{k},Y^{k})\}$ to train a local model $M_{k}$. The server $C$ aggregates the local model parameters $\{\theta_{k}\}$ from all participating clients to compute the global parameters $\theta_{g}=\sum_{k=1}^{K}\alpha_{k}\theta_{k}$, where $\alpha_{k}$ is the weight based on the size of each client’s local dataset, and $\alpha_{k}=\frac{|D^{k}|}{\sum_{k=1}^{K}|D^{k}|}$.
Under this setting, FedDDL constructs a structural causal graph to analyze the confounding factors involved in FL model inference. As shown in Figure 2, the main idea is to tackle the interference of client data heterogeneity ($S\rightarrow X$) and background ($B\rightarrow X$) factors in model decision-making, where $X$ is the input image. The intra-client deconfounding learning module to generate counterfactual data to improve data diversity, which can sever the link between background and specific class during inference (i.e., $B\nrightarrow X$). The inter-client debiasing learning module constructs casual prototypes $U_{c}=\{u_{c}^{1},...,u_{c}^{N}\}$ for each class based on object images, where $N$ is the number of classes. It uses $U_{c}$ to align heterogeneous representations (i.e., $f_{k}^{j}\mapsto u_{c}^{j},k=1,...,K,j=1,...,N$), which eliminates the impact of differences between models from diverse clients during aggregation (i.e., $S\nrightarrow X$).
@@ 4 The Proposed FedDDL Method
In this section, we describe the proposed FedDDL method in detail with reference to the schematic framework as shown in Figure 3.
@@ 4.2 Intra-Client Deconfounding Learning (DEC)
To address the issue of shortcut learning in which client-specific backgrounds misleading the model to form an association between the background and specific labels, the DEC module aims to sever the causal path from the background to the label during inference to encourage the model to focus on true causal relationships. To achieve this, it performs backdoor intervention to adjust the sample distribution $do(X)$. By applying the law of total probability, the inference involves both the direct causal impact of $X$ on $Y$ and the correlation confounded by $B$:
@@
which removes the path from $B$ to $X$ in Figure 2, thereby facilitating the model to approximate the causal intervention $P(Y\mid\text{do}(X))$ instead of the spurious correlation $P(Y\mid X)$. Specifically, the DEC module begins by decoupling images into objects and backgrounds. Then, it modifies the background factors associated with the objects. Inspired by the pre-trained model Grounding DINO Liu et al. (2025), the approach leverages text prompts $T$ to detect the corresponding targets:
@@ 5.1 Experiment Settings
5.1.1 Datasets
Following previous work on OOD generalization Qi et al. (2024a); Wang et al. (2022), experiments are conducted on two datasets: NICO-Animal and NICO-Vehicle Wang and et al. (2021). The statistical information and partitioning method of the datasets can be found in Table 1.
@@ 5.1.1 Datasets
5.1.2 Evaluation Metrics
Following previous studies McMahan and et al. (2017); Li et al. (2021), we evaluate the performance of all methods using Top-1 Accuracy, defined by the following equation:
$\text{Accuracy}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}(\hat{y}_{i}=y_{i}).$ (14)
$N$ represents the total number of samples. $\hat{y}_{i}$ is the predicted label for the $i$-th sample. $y_{i}$ is the true label for the $i$-th sample. $\mathbb{I}(\hat{y}_{i}=y_{i})$ is an indicator function, which is 1 if the predicted label equals the true label, and 0 otherwise.
5.1.3 Network Architecture
To ensure fair comparison, all methods adopt the same network architecture. In FedDDL, a unified architecture is used to process both raw images and those processed with Grounding DINO. Following prior studies Li and et al. (2021); Liu and et al. (2021), ResNet-18 He and et al. (2016) is selected as the backbone for both datasets.
5.1.4 Implementation Details
In the experiments, we set the local training epochs to 10 per global round for both datasets. The total number of communication rounds is 50, with 7 clients for both datasets. We used a client sampling fraction of 1.0 and employed SGD as the optimizer. During local training, the weight decay is set to 0.01, the batch size is 64, and the initial learning rate is 0.01 for both datasets. $\lambda$ is chosen from the set $\{0.1,1.0,2.0\}$. $\tau$ is selected from $\{0.5,0.07\}$. $\eta$ is tuned from $\{1,3,5\}$. The BOX_THRESHOLD and TEXT_THRESHOLD in the DINO model have been set to 0.3. For other methods, hyperparameters are tuned based on the corresponding papers.
@@ 5.2 Performance Comparison
This section presents a comparison of the FedDDL method with nine state-of-the-art (SOTA) methods, including FedAvg McMahan and et al. (2017), FedProx Li (2020), MOON Li and et al. (2021), FPL Huang and et al. (2023), FedIIR Guo and et al. (2023), FedDecorr Shi et al. (2024), FedHeal Chen et al. (2024), MCGDM Wei and Han (2024), and FedCCRL Wang and Tang (2024). The results derived from Table 2 are summarized below.
@@ 5.3 Ablation Study
  • By combining both modules, the best performance across all cases is achieved, highlighting the necessity of mitigating the interference caused by the two confounding factors on model inference. Notably, this provides a feasible solution for future research, offering an alternative to pursuing a singular approach.
@@ 5.5.2 Comparison of Visual Attention
This section analyzes the effectiveness of the deconfounding learning in local 1. We randomly selected test samples and visualized the visual attention of different methods using GradCAM Selvaraju et al. (2017); Ślazyk et al. (2022). As illustrated in Figure 8, FedAvg often focuses on the background of out-of-distribution samples, leading to classification errors due to its limited generalization ability. For example, in the case of a cat on grass, FedAvg may assign the label based on “grass,” leading to incorrect identification as a “dog,” since the training data only contains dogs on grass. Furthermore, FedDDL effectively focuses on the target object in test samples, enabling the model to eliminate background interference and make accurate predictions. This significantly enhances the generalization ability of local models and demonstrates that improving the performance of individual models also contributes to better collaboration.

### v2

@@ 1 Introduction
Federated out-of-distribution (OOD) generalization typically leverages data with diverse attributes across clients for collaborative modeling. The aim is to enhance model performance on unseen distributions Qi and et al. (2024); Qi et al. (2025b); Wang et al. (2024); Fu et al. (2025b). It allows FL clients to train models locally and iteratively exchange parameters with the server, enabling global aggregation without exposing private data Liao et al. (2024); Fu et al. (2025a); Hu et al. (2023); Yi et al. (2023). Despite its contributions to data privacy protection, existing studies show that attribute skew among sources often leads to performance degradation in FL models Fu et al. (2025a); Zhang and et al. (2024); Zhang et al. (2025a); Ren et al. (2025). This is primarily due to spurious correlations in local models, which hinder the formation of a robust global model during aggregation Yang et al. (2020); Hu et al. (2024); Cai et al. (2024b).
@@
Existing approaches for enhancing FL model performance on OOD samples can be broadly divided into two groups: 1) knowledge distillation-based methods Wei and Han (2024); Fan et al. (2025); Yu et al. (2023); Huang et al. (2025) and 2) data augmentation-based methods Chen et al. (2023); Shenaj et al. (2023); Park et al. (2024). The former either treats the teacher’s knowledge (e.g., the output of global or other local models) as a regularizer to provide additional supervision or decouples category-irrelevant features in the latent space, thereby mitigating the interference of irrelevant attributes. For instance, MCGDM Wei and Han (2024) shares classifiers across sources to match gradients both within and across domains to reduce overfitting to attribute features. DFL Luo et al. (2022) applies mutual information-based disentanglement to address negative transfer due to attribute skew. However, the suboptimal performance of the teacher model on OOD samples and the single environment limit the effects of knowledge distillation and attribute decoupling. The latter approach typically trains generators or uses pre-trained generative models to create data with diverse attributes (e.g., CCST Chen et al. (2023)). For example, FedCCRL Wang and Tang (2024) fuses different samples for data augmentation. However, they often introduce privacy risks with performance limited by the quality of the generated data.
@@ 2.1 Data Augmentation-based Methods
To mitigate attribute bias, data augmentation-based methods Wang and Tang (2024); Shenaj et al. (2023); Liu et al. (2021); Xu et al. (2023); Zhang et al. (2024); de Luca and et al. (2022); Zhang et al. (2025b); Qi et al. (2022) have been developed to improve model generalization to previously unseen attributes by boosting the diversity of the data attributes. These methods typically rely on two approaches: 1) training data generators to create novel samples, or 2) utilizing pre-trained diffusion models to enrich sample diversity. The first approach involves exchanging local information between clients to generate new samples from different domains. For instance, FIST Nguyen et al. (2024) and StableFDG Park et al. (2024) produce samples with varying styles by sharing style-related information across clients. The second approach leverages prompt-driven techniques to generate samples that align with certain criteria Morafah and et al. (2024); Zhao et al. (2023). Although enhancing diversity, these strategies often introduce privacy concerns. Moreover, the quality differences between generated and original samples also limit their effectiveness.
@@ 2.2 Knowledge Distillation-based Methods
Knowledge distillation-based approaches utilize generalized knowledge to guide local models in extracting shared features that are independent of specific attributes. These methods typically adopt two main strategies: 1) regularizing representations from different sources to align them, thus promoting consistency in representations across diverse contexts Wang et al. (2023); Qi et al. (2024, 2025a, 2023); Liu et al. (2021); Sun et al. (2023); Meng et al. (2024); Zhu et al. (2024); or 2) decoupling invariant features within the latent space to minimize the impact of confounding factors Yu et al. (2023); Yi et al. (2024); Yu and et al. (2011); Cai et al. (2024a). For instance, FedProc Mu et al. (2023) and FPL Huang et al. (2023) leverage prototypes to enforce consistent representation learning across all clients, thereby encouraging them to converge within a shared representation space. MCGDM Wei and Han (2024) mitigates overfitting within individual domains by employing both intra-domain and cross-domain gradient matching to improve generalization. Similarly, FedIIR Guo et al. (2023) enables the model to implicitly learn invariant relationships by capitalizing on prediction inconsistencies and gradient alignment across clients. Although these techniques have achieved notable performance improvements, they often rely on single-domain data, which hinders model transferability to OOD domains. Our FedDDL method bridges these gaps.
@@ 3 Preliminaries
An FL system typically involves multiple distinct clients, denoted as $\mathcal{S}=\{S_{1},S_{2},\dots,S_{K}\}$, and a central coordinating server $C$. Each $S_{k}$ utilizes its own private dataset $D^{k}=\{(X^{k},Y^{k})\}$ to train a local model $M_{k}$. The server $C$ aggregates the local model parameters $\{\theta_{k}\}$ from all participating clients to compute the global parameters $\theta_{g}=\sum_{k=1}^{K}\alpha_{k}\theta_{k}$, where $\alpha_{k}$ is the weight based on the size of each client’s local dataset, and $\alpha_{k}=\frac{|D^{k}|}{\sum_{k=1}^{K}|D^{k}|}$.
Under this setting, FedDDL constructs a structural causal graph to analyze the confounding factors involved in FL model inference. As shown in Figure 2, the main idea is to tackle the interference of client data heterogeneity ($S\rightarrow X$) and background ($B\rightarrow X$) factors in model decision-making, where $X$ is the input image. The intra-client deconfounding learning module to generate counterfactual data to improve data diversity, which can sever the link between background and specific class during inference (i.e., $B\nrightarrow X$). The inter-client debiasing module constructs class-wise causal prototypes $U_{c}={u_{c}^{1},\dots,u_{c}^{N}}$ from object images, and uses them to align heterogeneous representations across clients. This mitigates the influence of client-specific model differences during aggregation (i.e., $S\nrightarrow X$).
@@ 4 The Proposed FedDDL Method
This section presents the proposed FedDDL method with reference to the schematic framework shown in Figure 3.
@@ 4.2 Intra-Client Deconfounding Learning (DEC)
To address shortcut learning caused by client-specific backgrounds, the DEC module aims to block the causal link from background to label during inference, encouraging the model to focus on true causal relationships. To achieve this, it performs backdoor intervention to adjust the sample distribution $do(X)$. By applying the law of total probability, the inference involves both the direct causal impact of $X$ on $Y$ and the correlation confounded by $B$:
@@
which removes the path from $B$ to $X$ in Figure 2, thereby facilitating the model to approximate the causal intervention $P(Y\mid\text{do}(X))$ instead of the spurious correlation $P(Y\mid X)$. Specifically, the DEC module begins by decoupling images into objects and backgrounds. Then, it modifies the background factors associated with the objects. Inspired by the pre-trained model Grounding DINO Liu and et al. (2025), the approach leverages text prompts $T$ to detect targets:
@@ 5.1 Experiment Settings
Datasets
Following previous work on OOD generalization Qi and et al. (2024); Wang et al. (2022), experiments are conducted on two datasets: NICO-Animal and NICO-Vehicle Wang et al. (2021). Their statistics and partitioning method of the datasets can be found in Table 1.
@@
Evaluation Metrics
Following prior work McMahan et al. (2017); Li et al. (2021), we use Top-1 Accuracy to evaluate performance, which is defined as $\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}(\hat{y}_{i}=y_{i})$, where $N$ is the total number of samples, $\hat{y}_{i}$ is the predicted label, and $y_{i}$ is the ground-truth label.
Network Architecture
To ensure fair comparison, all methods adopt the same network architecture. In FedDDL, a unified architecture is used to process both raw images and those processed with Grounding DINO. Following prior studies moon; Liu et al. (2021), ResNet-18 He et al. (2016) is selected as the backbone for both datasets.
Implementation Details
We set the local training epochs to 10 per global round for both datasets. The total number of communication rounds is 50, with 7 clients for both datasets. We used a client sampling fraction of 1.0 and employed SGD as the optimizer. During local training, the weight decay is set to 0.01, the batch size is 64, and the initial learning rate is 0.01 for both datasets. $\lambda$ is chosen from the set $\{0.1,1.0,2.0\}$. $\tau$ is selected from $\{0.5,0.07\}$. $\eta$ is tuned from $\{1,3,5\}$. The BOX_THRESHOLD and TEXT_THRESHOLD in the DINO model have been set to 0.3. For other methods, hyperparameters are tuned based on the corresponding papers.
@@ 5.2 Performance Comparison
This section presents a comparison of the FedDDL method with nine state-of-the-art (SOTA) methods, including FedAvg McMahan et al. (2017), FedProx Li et al. (2020), MOON Li et al. (2021), FPL Huang et al. (2023), FedIIR Guo et al. (2023), FedDecorr Shi et al. (2024), FedHeal Chen et al. (2024), MCGDM Wei and Han (2024), and FedCCRL Wang and Tang (2024). The results derived from Table 2 are summarized below.
@@ 5.5.2 Comparison of Visual Attention
This section analyzes the effectiveness of the deconfounding learning in local 1. We randomly selected test samples and visualized the visual attention of different methods using GradCAM Selvaraju et al. (2017); Ślazyk et al. (2022); Lin et al. (2020); Meng et al. (2019). As illustrated in Figure 8, FedAvg often focuses on the background of out-of-distribution samples, leading to classification errors due to its limited generalization ability. For example, in the case of a cat on grass, FedAvg may assign the label based on “grass,” leading to incorrect identification as a “dog,” since the training data only contains dogs on grass. Furthermore, FedDDL effectively focuses on the target object in test samples, enabling the model to eliminate background interference and make accurate predictions. This significantly enhances the generalization ability of local models and demonstrates that improving the performance of individual models also contributes to better collaboration.
@@ 6 Conclusions and Future Work
Acknowledgments
This work is supported in part by the Joint Funds of the National Natural Science Foundation of China (Grant NO. U2336211); the Ministry of Education, Singapore, under its Academic Research Fund Tier 1; the RIE2025 Industry Alignment Fund – Industry Collaboration Projects (IAF-ICP) (Award I2301E0026), administered by A*STAR, as well as supported by Alibaba Group and NTU Singapore through Alibaba-NTU Global e-Sustainability CorpLab (ANGEL).